OpenStudy (anonymous):

3. d over three + 10 = 7 (1 point) 51 20 0 –9

OpenStudy (anonymous):

WILL GIVE MEDAL

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

for d/3 + 10 look for gcf w then the equation will be (d+30)/3 = 7 cross multiply 21= d+30 d=21-30 d= -9